# LiDAR Remote Sensing Meets Weak Supervision: Concepts, Methods, and Perspectives

### v1

@@ 2 LiDAR Interpretation with Weak Supervision
Weakly supervised learning in LiDAR remote sensing can be categorized into three types [47]: Incomplete Supervision, where only a subset of training data is labeled while the rest remains unlabeled; Inexact Supervision, where annotations are coarse-grained, such as providing scene-level labels instead of point-level labels in semantic segmentation; and Inaccurate Supervision, where annotation data may contain errors or noise. In addition, there is domain adaptation [48], which utilizes the fully labeled data in the source domain and the small amount or unlabeled data in the target domain to improve the task performance in the target domain by reducing the distribution differences between the two domains. Table 1 shows a summary of representative approaches.
@@ 2.1 Incomplete supervision
Incomplete supervision addresses the situation in which we are given a tiny portion of the labels as supervised signals and how to use the remaining vast amount of unlabeled data to ensure effective learning in the network [47, 49, 50], as shown in Fig. 3. It consists of two labeling strategies: parts of the point cloud scenes are densely labeled while the remainder are not labeled, and a small number of points are sparingly labeled in point cloud scenes. According to the relevant literature [47], the former is referred to as “semi-supervised learning” [47, 49, 50, 51], and the latter is referred to as “weakly supervised learning with sparse annotations” [52, 53, 54, 55].
@@ 2.1.1 Semi-supervised learning
It aims to improve network learning performance by utilizing a small portion of fully labeled training data from the scene and a large portion of unlabeled scene data [47, 49, 50, 51]. A lot of research has been done about the point cloud field, such as contrastive learning [56], consistency regularization [57], pseudo-label [58, 59],self-training [20], etc.
Semi-supervised 3D semantic segmentation is to completely utilize the data, with certain scenes fully labeled and the majority of the remaining scenes unlabeled, in order to achieve semantic annotation of all data. Li et al. [20] uses a self-training paradigm to improve the segmentation performance. [58, 59] generate pseudo-labels for unlabeled point clouds. Zhang et al. [57] utilizes consistent prediction of the teacher model and the student model to improve the performance of the segmentation network. Jiang et al. [56] apply contrast learning to 3D point cloud semantic segmentation. Some research [60, 56, 61, 62] applies semi-supervised learning for LiDAR semantic segmentation in autonomous driving. For instance, LaserMix [60] mixes laser beams from different LiDAR scans by exploiting a priori knowledge of spatial cues and maintains consistent prediction of the segmentation network before and after mixing.
In LiDAR remote sensing, most research focuses on airborne LiDAR point clouds. For instance, Liao et al. [30] propose a semi-supervised graph fusion framework to fuse spectral, spatial, and elevation features of LiDAR and hyperspectral data. Li et al. [31] propose a two-stage classification method with pseudo-label contrast learning to solve the time-consuming and laborious problem of jointly classifying labeled samples with HSI and LiDAR. Li et al. [32] generate high-confidence pseudo-labels using a novel label calibration module to enhance semi-supervised learning. Guo et al. [33] first extract fused features from all samples and then classify them using a limited number of labeled samples. Pu et al. [34] employ an unsupervised encoder-decoder structure combined with a data-specific branch to perform joint training, enabling the extraction of effective multimodal representations even with limited labeled samples. Additionally, Dersch et al. [35] utilize semi-supervised learning for multi-class tree crown delineation.
Multi-source data fusion is increasingly favored in LiDAR remote sensing for information complementarity. Combining hyperspectral and LiDAR data addresses their individual shortcomings: hyperspectral’s susceptibility to SODS/SSDO hinders accurate classification, while LiDAR lacks object detail. Fused data provides comprehensive, mutually reinforcing surface information, boosting classification accuracy and robustness significantly.
@@
Zhao et al. [64] propose a robust 3D object detector trained via the Mean Teacher paradigm, incorporating point cloud-specific perturbations and consistency loss. Similarly, Wang et al. [65] utilize a teacher-student mutual learning framework to propagate information from labeled data to unlabeled data in the form of pseudo-labels. Outdoor LiDAR point clouds present unique challenges such as complex structures [60] and low point density [66]. Research on 3D point cloud object detection in LiDAR remote sensing is limited. Current research [65, 67, 68, 69, 70, 71, 72, 73] is largely focused on LiDAR point clouds for autonomous driving. For instance, 3DIoUMatch [65] set a milestone in semi-supervised 3D object detection by creating a novel pseudo-label filtering method guided by both detection scores and IoU. Qi et al. [67] leverage temporal points through multi-frame object detection to achieve an offboard 3D object detection pipeline. Xu et al. [68] generate high-quality pseudo-labels through an adaptive class confidence selection module and utilize a holistic point cloud augmentation module to improve robustness. Wong [73] takes into account the non-coincidence and weak correlation of the target objects in the point cloud by introducing two mechanisms: strict thresholding and filter switching, to realize that only the truth-determining pseudo labels are retained and the other fuzzy labels are pruned.
@@
Compared to semi-supervised 3D semantic segmentation and object detection, research on semi-supervised 3D instance segmentation is limited [74, 75, 76, 77]. For instance, Liao et al. [74] force the predictions of different perturbations of the input point cloud to remain constant to achieve semi-supervised point cloud instance segmentation. Chu et al. [76] use pseudo semantic labels and pseudo offset vectors for semantic and instance-level supervision, with an object-level denoising module and a bidirectional mutual enhancement mechanism to iteratively improve pseudo label quality. Liu et al. [77] apply a contrast learning strategy to guide multi-stage network training.
@@ 2.1.2 Sparse-supervised learning
It refers to the scenario in 3D point cloud data where only a small number of points are sparsely labeled, and these limited annotations, combined with large amounts of unlabeled data, are used to train the network effectively. The research methods include consistency regularization [19, 78], contrastive learning [79], self-training [80], label propagation [52, 81], pseudo-label [36, 37, 81, 82], active learning [22, 80], and others.
@@
Sparse-supervised 3D semantic segmentation refers to performing semantic segmentation on 3D point cloud data where only a subset of points are labeled, while most remain unlabeled. The goal is to train a model capable of accurately segmenting the entire point cloud using limited labeled information.
@@
A large number of well-established works [19, 52, 78, 83, 22, 84, 85] have emerged. Xu et al. [19] employ a teacher-student model and leverage active learning to select and label the most informative samples, achieving good results with very little labeled data. Considering information redundancy in existing active learning strategies, Xu et al. [22] design a feature distance suppression module to select the important points to be labeled. Liu et al. [52] label one point for each object and propagate the labels to unknown regions through a graph propagation module. [78, 83] force predictive consistency of the original and perturbation branches. Hu et al. [84] leverage the semantic similarity of neighboring 3D point clouds by encoding raw point clouds and their neighbors into vectors to predict semantic labels. [86, 80, 81, 82, 87, 88] apply similar strategies like pseudo-labeling, active learning, and self-training to LiDAR point clouds in autonomous driving. For instance, Hu et al. [80] propose a new active learning strategy for 3D LiDAR semantic segmentation based on the inconsistency of inter-frame prediction to estimate model uncertainty and further ensemble self-training to further improve performance. Zhao et al. [81] propagate pseudo-labels between adjacent frames and select reliable labels for learning.
Currently, there is very limited research on LiDAR remote sensing. For airborne laser point cloud, Wang et al. [36] propose a pseudo-label-assisted point cloud segmentation method, which generates pseudo-labels by setting an adaptive threshold for the prediction results of unlabeled data, after which the pseudo-labels are used for iterative training. In addition, they penalize the uncertainty in the prediction class by introducing entropy regularization while imposing consistency constraints between the predictions and the integrated predictions to improve the robustness of the network [18]. For mobile laser point cloud, Lei et al. [37] propose a multi-branch weakly supervised learning network (WSPointNet), which contains an integrated prediction constraint branch to improve the stability of point cloud prediction, a contrast-guided entropy regularization branch to prevent model overfitting, and an adaptive pseudo-label learning branch. Subsequently, they propose a novel weakly supervised MLS point cloud semantic segmentation method by incorporating data augmentation and active learning strategies into WSPointNet [38].
@@
Sparse-supervised 3D object detection refers to detecting objects in 3D point clouds, where only a subset of objects are labeled while most remain unlabeled. The task involves predicting the object categories and their corresponding 3D bounding boxes. Most research is focused on LiDAR point clouds for autonomous driving, with fewer studies targeting indoor point clouds and virtually none addressing LiDAR remote sensing.
Xu et al. [89] implement a weakly supervised object detection method for indoor point clouds by using synthesized 3D shapes to convert weak labels into fully annotated virtual scenes as stronger supervision. For LiDAR Point Cloud in autonomous driving, Liu et al. [54] mine positive instances through a missing-annotated instance mining module and design a reliable background mining module and a point cloud filling data augmentation strategy to generate confidence data for iterative learning. Meng et al. [90] utilize a small number of weakly annotated scenarios to generate cylindrical object proposals and refine cylindrical proposals using a small number of accurately labeled object instances to obtain cuboids and confidence scores. Xia et al. [91] enhance feature recognition by introducing a multi-class contrast learning module (MCcont) and supervise detector training with the introduction of an instance feature mining module and a mark-to-pseudo contrast learning module. Considering that existing weakly supervised and semi-supervised methods are difficult to adapt to different scenarios (e.g., indoor or outdoor), Zhang et al. [92] utilize a small amount of full box-level annotations and a large amount of point-level annotations to generate high-quality pseudo-boundary boxes for training any 3D detector by modeling the global interactions between the LiDAR points and the corresponding weak labels. Yang et al. [93] utilize a large number of coarse cluster-level labels to learn semantics and a number of expensive box-level labels to learn accurate poses and shapes. Gao et al. [94] improve PointDETR [95] and propose PointDETR3D, a faculty-student framework for weakly supervised 3D detection.
@@
Sparse-supervised 3D instance segmentation refers to performing instance-level segmentation on 3D point cloud data where only a subset of points or instances are labeled. The goal is to assign each point to a specific instance ID while predicting the instance’s category label.
Hou et al. [96] observe that PointContrast [97] uses a simple pretext task based on contrast learning that focuses only on point-level correspondence matching and propose a method to integrate spatial information into the contrast learning frame. Tao et al. [98] transform instance location annotations into segment-level labels via oversegmentation and then generate point-level pseudo-labels from these using a segment grouping network (SegGroup). Tang et al. [99] extend point-level labels to the superpoint level and leverage inter-point affinities to generate high-quality pseudo-labels. Liu et al. [100] annotate one point for each instance and further devise a new similarity metric for clustering to generate pseudo-labels. Dong et al. [101] propose a weakly supervised 3D instance segmentation method that only requires categorical semantic labels and not instance-level labels as supervision. They propagate semantic and instance information to unknown regions by self-attention and cross-graph random wandering, respectively [102].

### v2

@@ 3 LiDAR Interpretation with Weak Supervision
WSL in LiDAR remote sensing can be categorized into three types [105]: Incomplete Supervision, where only a subset of training data is labeled while the rest remains unlabeled; Inexact Supervision, where annotations are coarse-grained, such as providing scene-level labels instead of point-level labels in semantic segmentation; and Inaccurate Supervision, where annotation data may contain errors or noise. In addition, there is domain adaptation (DA) [106], which utilizes the fully labeled data in the source domain and the small amount or unlabeled data in the target domain to improve the task performance in the target domain by reducing the distribution differences between the two domains. Furthermore, domain generalization (DG) aims to learn a model from data in one or more related source domains, enabling it to generalize effectively to any unseen target domain [107]. Because research on object detection and instance segmentation in LiDAR remote sensing is limited, existing work has focused primarily on semantic segmentation. Therefore, Table 1 lists commonly used semantic segmentation datasets, while Tables 2 and 3 summarize representative LiDAR remote sensing methods and their accuracies on the evaluated datasets, respectively.
@@ 3.1 Incomplete supervision
Incomplete supervision refers to the scenario where, with only limited labeled data providing supervised signals, a large amount of unlabeled data is leveraged to ensure effective network learning  [105, 108, 29]. Fig. 3 provides a schematic illustration redrawn from [29]. Drawing from the categorization in [105, 29], we divide incomplete supervision into four main techniques: (i) Learning strategies that densely annotate partial regions of point cloud scenes while leaving the rest unlabeled, commonly known as semi-supervised learning  [105, 108, 29]; (ii) Strategies that sparsely annotate only a few points in point cloud scenes, termed WSL with sparse annotations  [62]; (iii) Active learning methods in which the model proactively selects the “most informative” samples from unlabeled data for human annotation; (iv) Few-shot learning that utilizes abundant labeled samples from base classes and limited samples from novel classes to train the model, thereby reducing annotation costs for novel classes.
@@ 3.1.1 Semi-supervised learning
It aims to improve network learning performance by utilizing a small portion of fully labeled training data from the scene and a large portion of unlabeled scene data [105]. Extensive research has been conducted in the point cloud field, such as contrastive learning [109], consistency regularization [110], pseudo-labeling [111], self-training [22], etc.
Semi-supervised 3D semantic segmentation aims to fully leverage all data in a setting where a few scenes are fully labeled while the majority remain unlabeled, in order to achieve comprehensive semantic annotation. Li et al. [22] introduce a GAN-like discriminator within a self-training framework to evaluate pseudo-label confidence. This approach proves more reliable than solely relying on model outputs, but the adversarial training can introduce instability. Zhang et al. [110] employ the classic Mean Teacher model, utilizing unlabeled data through consistency regularization; however, this class of methods essentially treats a point cloud as a collection of independent points and fails to fully exploit its rich geometric structure priors. Recognizing the importance of structural information, subsequent research gives rise to methods centered on “superpoints”. Works by Cheng et al. [111] and Deng et al. [112] first over-segment point clouds into superpoints with local consistency, elevating the segmentation task from the point to the region level to better utilize local geometric and semantic consistency. Further research applies semi-supervised methods to LiDAR semantic segmentation in autonomous driving scenarios. For example, LaserMix [113] leverages the strong spatial priors of LiDAR data by mixing specific laser beams from different scan frames to generate new samples while maintaining prediction consistency, achieving significant results. However, its heavy reliance on sensor and scene priors limits its generalizability to unstructured indoor scenes or general 3D models.
In LiDAR remote sensing, most research focuses on ALS point clouds. Early work by Liao et al. [55] proposes a semi-supervised graph-fusion framework that combines spectral, spatial, and elevation features from LiDAR and hyperspectral data. With the advent of deep learning, attention shifts to more sophisticated neural architectures. Li et al. [6] introduce a ternary deep network (TSDN) that uses dedicated branches to address modality heterogeneity and adopts a multi-model consensus strategy to purify pseudo-labels, thereby enhancing the robustness of semi-supervised learning. Guo et al. [57] adopt a two-stage “unsupervised pretraining—supervised fine-tuning” paradigm, learning general features via reconstruction. Pu et al. [58] employ an unsupervised encoder–decoder with data-specific branches for joint training, enabling effective multimodal representations from limited labeled samples. Additionally, Dersch et al. [59] apply semi-supervised methods to multi-class tree-crown delineation. More recently, Li et al. [56] adopt contrastive learning, replacing pixel-level reconstruction objectives with feature-space constraints that enforce intra-class compactness and inter-class separability to obtain more discriminative representations. To obtain more discriminative representations. Notably, many of these approaches leverage multi-source data fusion, combining LiDAR’s geometric strengths with hyperspectral data to overcome issues like same object different spectra (SODS)/same spectrum different objects (SSDO). This fusion, while powerful, increases annotation costs, further reinforcing the necessity of semi-supervised techniques to make such advanced analysis practical and scalable.
@@
Zhao et al. [115] propose a robust 3D object detector trained via the Mean Teacher paradigm, incorporating point cloud-specific perturbations and consistency loss. Similarly, Wang et al. [116] utilize a teacher-student mutual learning framework to propagate information from labeled data to unlabeled data in the form of pseudo-labels. Outdoor LiDAR point clouds present unique challenges such as complex structures and low point density [117]. Research on 3D point cloud object detection in LiDAR remote sensing is limited. Current research is largely focused on LiDAR point clouds for autonomous driving [118, 119]. For instance, 3DIoUMatch [116] introduces pseudo-label filtering based on IoU prediction to eliminate low-quality annotations. However, the reliability of this filtering is compromised in sparse or occluded scenes, where noise in the IoU predictions undermines the selection process. Qi et al. [120] leverage full point cloud sequences offline to generate high-quality “auto-labels” significantly enhancing label precision. Nevertheless, its offline nature and substantial computational cost render it unsuitable for real-time deployment. Most recently, SSF3D [121] adopts a “strict semi-supervised” strategy using stringent thresholds and filter switching to ensure pseudo-label quality, but this risks biasing the model towards easy samples and thus impairing its generalization ability on more challenging instances.
@@
Compared to semi-supervised 3D semantic segmentation and object detection, research on semi-supervised 3D instance segmentation is limited [122, 123]. For instance, Liao et al. [124] leverage consistency regularization—enforcing invariant predictions under input perturbations—to exploit unlabeled data, yet the shallow self-supervised signal struggles to capture the deep correlation between instances and semantics, offering limited improvement when annotations are scarce. Chu et al. [123] propose a self-training framework that facilitates mutual reinforcement between semantic and instance pseudo-labels and introduces a learnable “proposal re-correction module” to denoise these pseudo-labels at the object level, significantly enhancing pseudo-label quality. However, the two pseudo-label types are prone to error coupling, and the framework’s effectiveness depends on the re-correction module’s performance. Liu et al. [125], in contrast, construct regional units via over-segmentation and simultaneously model semantic and instance boundaries using a boundary-aware energy function and contrastive learning. This approach introduces a stronger structural prior, but its performance is constrained by the initial over-segmentation quality, and the joint optimization of the energy function and contrastive objective presents additional challenges.
@@ 3.1.2 Sparse-supervised learning
It refers to the scenario in 3D point cloud data where only a small number of points are sparsely labeled, and these limited annotations, combined with large amounts of unlabeled data, are used to train the network effectively. The research methods include consistency regularization [21], contrastive learning [126], self-training [127], label propagation [128], pseudo-labeling [64], active learning [129], and others.
@@
Sparse-supervised 3D semantic segmentation refers to performing semantic segmentation on 3D point cloud data where only a subset of points is labeled, while most remain unlabeled. The goal is to train a model capable of accurately segmenting the entire point cloud using limited labeled information.
@@
A large number of well-established works have emerged [21, 130]. Xu et al. [21] employ a teacher-student model and leverage active learning to select and label the most informative samples, achieving good results with very little labeled data. Considering information redundancy in existing active learning strategies, Xu et al. [129] design a feature distance suppression module to select the important points to be labeled.
Hu et al. [63] tackle the challenge of large-scale point cloud segmentation at a minimal cost by querying local semantic information to efficiently propagate sparse supervisory signals, although the core assumption of local semantic homogeneity compromises the segmentation accuracy at object boundaries. Some research applies similar strategies like pseudo-labeling, active learning, and self-training to LiDAR point clouds in autonomous driving [127, 128, 131]. For instance, Hu et al. [127] propose a new active learning strategy for 3D LiDAR semantic segmentation based on the inconsistency of inter-frame prediction to estimate model uncertainty and further ensemble self-training to improve performance. Zhao et al. [128] bidirectionally propagate and refine pseudo-labels between sparsely annotated full frames to address the high annotation cost in the semantic segmentation of sequential LiDAR point clouds. However, the segmentation performance relies heavily on the accuracy of the scene flow estimation and degrades as the interval between annotated frames increases due to error accumulation.
Currently, extensive research has been conducted in LiDAR remote sensing, and studies addressing point clouds acquired by different platforms mainly include the following: for ALS point clouds, Wang et al. [60] introduce an adaptive-threshold strategy to select reliable pseudo-labels; moreover, they penalize prediction uncertainty with an entropy regularizer and impose consistency between single and fused predictions to improve network robustness [20]. Zhang et al. [61] propose the Perturbed Self-Distillation (PSD) framework, which builds perturbed branches to capture point-cloud graph topology and enforce prediction consistency for information transfer between labeled and unlabeled points; they also explore colorization-based self-supervised pretraining to transfer priors from unlabeled data into weakly supervised networks [62]. The Semantic Query Network (SQN) [63] exploits local semantic homogeneity and aggregates multi-scale contextual features (e.g., via interpolation) to implicitly strengthen supervision under extremely sparse annotations (e.g., 0.1%). Extensions of the teacher–student paradigm, such as HSCN [70], DCUF-Net [73] and DSDCL [72], further enhance consistency learning and discriminative feature learning by introducing multi-scale geometric similarity losses, data-and-feature dual perturbations with uncertainty-aware mechanisms, and feature-space pull/push strategies, respectively; DR-Net [71] designs a Dilated Region Feature Aggregation (DRFA) module to aggregate local features at multiple dilation scales and enrich representations.
For MLS point clouds, Guinard et al. [132] manually select representative points and partition the cloud into geometrically consistent regions, applying a conditional random field (CRF) for context-aware classification to achieve accurate urban scene labeling under severely limited supervision. Lei et al. [64] propose the multi-branch weakly supervised WSPointNet, which introduces a temporal dimension and maintains an exponential moving average (EMA) of per-point predictions to form a more stable “ensemble teacher” that the consistency objective aligns to, thereby improving robustness. MSCs [65] impose scene-level constraints at encoder, decoder and classifier stages to mitigate class imbalance, and Chen et al. [66] exploit hand-defined line and plane points as extra supervision to guide feature learning in a line–plane–object hierarchy.
For multispectral (MS-LiDAR) point clouds, methods aim to integrate multimodal information efficiently and reduce reliance on full supervision: FPWS-Net [67] uses main/auxiliary semantic inference modules constrained by consistency and mutual pseudo-label losses to enforce agreement on unlabeled data, while SCSQ-Net [68] employs a shared kernel point convolution (SKPC) backbone that shares weights across encoding and semantic inference stages to integrate neighborhood information efficiently. To address long-tailed class imbalance in TLS datasets, Li et al. [69] decouple training into two alternating stages: a feature-learning stage that uses resampling or reweighting to force the feature extractor to learn all classes more evenly (especially rare tail classes), and a classifier fine-tuning stage in which the extractor is frozen and the classifier is adjusted using the original biased data, thereby balancing discriminative power and distributional bias. Overall, weakly supervised methods across point-cloud types—through targeted mechanisms such as pseudo-label filtering, self-distillation, temporal smoothing, multi-scale consistency, and task decoupling—substantially reduce annotation cost while improving generalization and robustness in real-world scenarios.
@@
Sparse-supervised 3D object detection refers to detecting objects in 3D point clouds, where only a subset of objects is labeled while most remain unlabeled. The task involves predicting the object categories and their corresponding 3D bounding boxes. Most research is focused on LiDAR point clouds for autonomous driving, with fewer studies targeting indoor point clouds, while its application in LiDAR remote sensing remains significantly underexplored.
Meng et al. [133] pioneer the use of bird’s-eye-view center-point clicks as extremely weak labels to generate cylindrical candidate boxes, which are then optimized with a few precise annotations. Although this approach significantly reduces the annotation burden, it relies on the quality of the initial candidates, exhibits limited generalization to objects with complex or highly variable shapes, and still requires precise annotations. To compensate for the geometric information missing in weak labels, Xu et al. [134] utilize a 3D shape library to construct virtual scenes and transfer knowledge via domain adaptation. However, this method is constrained by the domain gap between virtual and real-world data and the availability of high-quality shape libraries. For LiDAR point clouds in autonomous driving scenarios, Liu et al. [135] discover positive samples through a missing-label instance mining module, design a reliable background mining module, and propose a point cloud completion data augmentation strategy to generate high-confidence data for iterative learning. Zhang et al. [136] leverage a small number of fully-annotated box labels and a large number of weakly-annotated point labels. By modeling the global interaction between LiDAR points and weak annotations, they generate high-quality pseudo bounding boxes to train any 3D detector.
@@
Sparse-supervised 3D instance segmentation refers to performing instance-level segmentation on 3D point cloud data where only a subset of points or instances is labeled. The goal is to assign each point to a specific instance ID while predicting the instance’s category label.
PointContrast [137] pioneers scene-level unsupervised pre-training using contrastive learning, but its focus on point-level consistency neglects global spatial structure, limiting its performance on complex tasks such as instance segmentation. SegGroup [138] weakens supervision to a single-point annotation per instance and generates pseudo-labels through segment grouping, which significantly reduces annotation costs but is highly dependent on the initial over-segmentation and grouping strategy, making it prone to error propagation. ClickSeg [139] refines single-point supervision to the “click-level”, using annotated points as initial seeds for k-means. While this simplifies the pipeline, it remains constrained by the inherent resolution limitations of clustering in dense or occluded scenes. Dong et al. [140] advance this line of research further by relying solely on semantic labels and learning an “objectness” signal to separate instances. Despite the significance of this step, the core challenge remains how to reliably generate initial instance seeds in regions with clustered same-class objects to avoid erroneous merging or over-segmentation.
3.1.3 Active learning
Active learning addresses scenarios with abundant unlabeled data and scarce labeled data. It uses an intelligent query process to selectively sample the most informative instances for expert annotation, aiming to maximize model performance at a minimal annotation cost. Active learning has achieved great success in 2D images. Its methods can be categorized into three main types: uncertainty-based approaches [141], which select samples with low predictive confidence to maximize information gain and refine decision boundaries; diversity-based approaches [142], which prioritize representative and diverse samples to ensure comprehensive coverage of the data distribution and mitigate redundancy inherent in uncertainty methods; and hybrid approaches [143], which balance informativeness and representativeness, typically by first filtering candidates based on uncertainty and then removing redundancy through diversity criteria (or vice versa).
Early works (e.g., ReDAL [144] and Shi et al. [145]) pioneer shifting the active-learning selection unit from whole scenes to subregions (regions/super-points), using measures such as entropy, color, and structural complexity to identify informative areas while promoting inter-region diversity to reduce redundancy. However, these methods rely on pre-segmentation, which can introduce partitioning bias and leave substantial redundancy inside selected regions. To address this, Annotator [146] proposes a voxel-centric online selection strategy: a “voxel confusion degree” (VCD) that jointly captures uncertainty and diversity, and source-domain pretraining to mitigate cold-start issues, markedly improving labeling efficiency in large-scale scenes. Nearly concurrently, Xu et al. [129] abandon region-based selection and introduce a hierarchical point-wise scheme that uses a multi-scale uncertainty module (HMMU) to assess point-level value and a feature-distance suppression (FDS) mechanism to ensure representativeness and dispersion—enabling more precise selection under very low annotation budgets. Despite these advances, manual annotation costs and potential annotator bias remain, and most methods focus on segmentation; their extension to other 3D tasks warrants further study.
Recent work in LiDAR remote sensing has advanced active-learning methods to reduce annotation cost. Early work by Polewski et al. [74] combines active learning and semi-supervised learning within a Rényi-entropy regularization framework: by minimizing the entropy of unlabeled samples, it selects the most informative samples for annotation and substantially improves classification accuracy in early training. Luo et al. [75] adopt an active-learning scheme that iteratively labels a small number of unlabeled supervoxels using a neighborhood-consistency prior. With the rise of deep learning for point clouds, Lin et al. [76] integrate active learning with deep networks (e.g., PointNet++) and propose a more efficient block-level segmentation entropy for uncertainty estimation. This advance improved feature learning and model performance but revealed a core limitation: uncertainty-based selection tends to pick “difficult” samples that are not necessarily representative, leading to spatially and structurally redundant (low-diversity) selections. To mitigate this, Lin et al. [77] later introduce an optimized method that combines point entropy, segmentation entropy, and mutual information to guide sample selection, and adopt incremental fine-tuning to reduce annotation needs and training time. Kölle et al. [78] combine active learning with crowdsourced annotation, using a Random Forest to select a small set of points from ALS point clouds and textured 3D meshes for efficient coupled semantic segmentation. SSDR-AL [79] uses a graph-reasoning network over superpoint diversity to progressively select informative and representative samples and then applies a noise-aware iterative annotation strategy. Kölle et al. [80] further shift from a model-centric to a data-centric paradigm and validate active-learning performance. More recent studies by Wang et al. [81] and Dai et al. [82] extend these ideas to extremely sparse supervision (One-Class-One-Click) and propose more efficient uncertainty measures (e.g., “Temporal Output Discrepancy”) that require no extra model ensembling, as well as human-inspired Spatial Incremental Sampling strategies. These advances substantially improve label efficiency but introduce new challenges—additional hyperparameters and open questions about defining and measuring diversity. Finally, DAAL-WS [83] employs an ensemble-prediction-based active-learning policy (PLEP) that selects points by point-level uncertainty to achieve high-accuracy 3D scene understanding with very few labeled samples.
3.1.4 Few-shot learning
Few-shot learning aims to rapidly adapt to novel classes and achieve strong generalization when only a very small number of labeled examples are available per class in the target domain. This paradigm has made significant progress in 2D images. Current mainstream methods typically rely on the meta-learning framework and can be further categorized into three types: metric-based methods [147], which learn an appropriate similarity metric space to perform classification predictions using the similarity (such as cosine similarity or Euclidean distance) between query samples and support set samples; optimization-based methods [148], which optimize the model’s initialization parameters or optimization process via meta-learning, enabling rapid adaptation to new tasks with few samples through a limited number of gradient steps; and augmentation-based methods [149], which generally enable rapid adaptation through internal model mechanisms or data/memory augmentation.
As a pioneering work, AttMPTI [150] successfully transplants the prototype-learning paradigm from 2D images to point clouds — using a pretrained encoder to distill class prototypes from few support samples to guide segmentation. However, this direct transfer reveals a fundamental weakness: large scene- and object-shape discrepancies between the support and query sets make prototypes distilled from isolated support examples fragile and unrepresentative in complex query scenes. To bridge this gap, both BFG [151] and 2CBR [152] explore a “pretraining + prototype adaptation” framework, introducing query information to dynamically adjust or “correct” the prototypes. Although these methods yield some performance gains, they do not address deeper limitations such as domain bias introduced by the pretrained encoder and the high training cost; consequently, model efficiency and generalization remain fundamentally constrained. A true paradigm shift appears in works represented by TaylorSeg [153], which abandon costly pretraining in favor of training-free geometric fitting or lightweight learnable modules — greatly improving efficiency and avoiding domain bias at its root, while potentially sacrificing the feature richness afforded by large-scale pretraining. Meanwhile, COSeg [154] takes a more critical view of the task formulation itself, exposing issues in earlier benchmarks such as “foreground leakage” that fundamentally undermine the reliability of some reported metrics.
In LiDAR remote sensing, few-shot learning remains underexplored. Zhang et al. [86]’s Pole-NN achieves a breakthrough in few-shot classification of pole-like objects using a training-free, non-parametric feature extractor, but it is restricted to specific object classes and does not readily scale to complex urban scene segmentation. Huang et al. [85] reduce annotation requirements for full-scene segmentation by combining sparse annotations with distribution calibration; however, this pretraining-dependent approach struggles to generalize under large domain shifts (e.g., different sensor types). To tackle cross-domain transfer, Dai et al. [84] propose Thr-MPRNet, a meta-learning multi-prototype relation network that transfers knowledge from photogrammetric point clouds to airborne laser scanning (ALS) data, yet meta-learning methods are prone to catastrophic forgetting when novel classes appear continuously in dynamic scenes. Addressing this, Dai et al. [87] introduce a cross-domain incremental feature-learning framework that uses a distillation module to dynamically adapt to newly emerging target-domain classes while retaining transferred knowledge. Most prior methods, however, treat geometric, intensity and other cues uniformly and thus ignore modality heterogeneity; Yuan et al. [88] explicitly point out this shortcoming and propose a multi-branch feature extractor that separates geometric and reflectance features to improve transfer and incremental learning in complex ALS scenarios. Together, these developments signal a shift from purely algorithmic strategies toward more fine-grained feature-representation learning.
3.1.5 Discussions for Incomplete supervision
Unlike optical imagery, LiDAR point clouds are inherently sparse, irregularly sampled, and spatially discontinuous, especially in regions affected by occlusion or variable scanning patterns. Such modality characteristics exacerbate incomplete supervision, where labeled points cover only a small fraction of the 3D scene and exhibit strong spatial and structural imbalance. The core strategy of weakly supervised learning under incomplete labels is to propagate limited annotations through pseudo-labeling and consistency regularization, enabling the model to leverage the abundant unlabeled data. However, applying these methods directly to LiDAR faces modality-specific challenges: sparse and uneven sampling interrupts label propagation, vertical structural layering leads to semantic misalignment between canopy and ground points, and noise distribution varies drastically across scanning patterns.
To address these issues, LiDAR-specific adaptations primarily introduce structure-aware pseudo-label generation and geometry-constrained consistency learning. Specifically, label propagation relies not only on feature similarity but also on geometric primitives (e.g., planarity, linear structures, and surface normals) for guidance, ensuring that pseudo-labels conform to physical structures. Multi-scale neighborhood queries are employed to better exploit the local and global geometry of point clouds, while customized data augmentation strategies enable learning representations robust to point sparsity, scanning variations, and noise. These targeted strategies mitigate the inherent sparsity, structural complexity, and noise of LiDAR, enabling weakly supervised methods to reliably amplify sparse annotations, narrow the supervision gap, and preserve geometric fidelity in large-scale 3D environments.